OpenStudy (anonymous):

Please help. Long problem. Will give medal and fan whoever succeeds in assisting me. Thank you.

OpenStudy (anonymous):

Follow the 5 steps below to complete this problem. (4 points) Step 1: Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: Drive a boat Drive a jet ski Step 2: Select a current speed of the water in mph. Step 3: Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current). Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (anonymous):

so first pick a name and do u want them on the boat or ski?

OpenStudy (anonymous):

Carla. Boat.

OpenStudy (anonymous):

so Carla pedals his paddle boat up the stream to visit bob in 2 hrs After his visit, he heads down stream, and was glad to be be home in 1 hr. the speed of the current was 2 mph, : Let s = his pedal speed in still water : write a dist equation; dist = time * speed : Upstream dist = downstream dist 2(s-2) = 1(s+2) 2s - 4 = s + 2 2s -s = 2 + 4 s = 6 mph is his speed in still water : : Confirm the solution by finding the distance, they should be equal 2(6-2) = 8 mi 1(6+2) = 8 mi

OpenStudy (anonymous):

or drive the boat not paddle

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

hmmmm I think so

OpenStudy (anonymous):

Alright, thank you very much.